Two systems of equations are given below.
For each system, choose the best description of its solution. If applicable, give the solution.
-x+2y=8
x-2y=8
and
x-2y=-4
-x+2y=4

Answer by ewatrrr(24785)


Hi
-x+2y=8
x-2y=8 |solving by first eliminating y (adding the two equatons)
0 ≠ 16 There is NO solution to this system
x-2y=-4
-x+2y=4 |solving by first eliminating y (adding the two equatons)
0 = 0 these are the same lines, infinite points in common for this system